Shift Summary Day 23 Dec 1999

Douglas A. Jensen (djensen@fnal.gov)
Thu, 23 Dec 1999 15:40:30 -0600 (CST)

Shift Summary for Day shift, 23 Dec 99

On Shift: Viv, DougJ

Hours of beam available at all intensities: 6.0/8.5
Hours of beam available at nominal intensity: 5.9
Hours of E799 data (special runs not included): 5.9
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift:
Run 15272 - 16.4M events - BEAM_799

Tapes brought to FCC this shift: none
========================================================================
Accesses during shift: one, Bob T, Julie, JasonH, Ken, SashaV
during beam off time.

========================================================================
Overview of shift: E799 datataking, except for beam off from 13:00 to
end of shift.

CsI gains continue changing while the blockhouse being heated by
blockhouse heater. (CsI-top = 18.6C last shift, 18.9C this shift
at the end of shift)

Otherwise, smooth running.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15272 07:00 new shift, dj dribbles in 5 min late.
07:55 Mount onsplit tape UPm048 (needed more labled).
08:03 " " UPh045.
09:40 MCR says tey will tune for a bit more intensity.
currently at about 8.5, 9.5 requested. BAO 1.04.
10:07 Channel 1096 continues to be destroyed...
(known problem.. )
10:15 beam a bit over requested.
11:00 missed 4 of last 9 spills.
11:40 10 min off later - booster tweeking. MCR called,
might go to an hour. Then, .1 - 1 hrd est at 1pm
Missed 6 spills.
13:00 Beam off for MCR purposes. Lots of people on
KTeV access.
BobT tweeked a fan
Julie tweeked a base
Jason tweeked a cable - see log for details
15272 13:00 End Run - tapes well along toward filling.

15275 14:10 Start run, in expectation of beam

14:45 At 14:00, the MCR est was beam at 14:45 -
Booster water leak.

15275 40 15:03 Beam! 3.4E12 this spill
a few lousy spills, Booster problems
then KTeVANA errors, pipeline resetes....
60 15:30 Real spills start - finally
========================================================================
Follow up Notes on Previous Shift
- Taking as much good data as possible
- Watched channel 968. It has a 6 count ped shift now.

========================================================================
Items to Follow up on Next Shift

- Take as much good data as possible.
=======================================================================
We are now looking to Santa Clause to solve the "KTeV Xerography Problem"!
... any chance he will come early? any chance he will come to
KTeV? ( there are clearly those in the high regions of management
who must not feel we are very deserving, as we keep loosing beam
time. Does that mean we won't get a solution to the
"KTeV Xerography Problem" either? )
=======================================================================

*** END of REPORT ***